1) What is the wavelength of radiation with a frequency of 7.30 x 10^15 sec^-1?
2) Find the frequency of radiation that has energy of 3.8 x 10^-10 J.
1
Expert's answer
2019-11-14T06:17:07-0500

Q98649


Solution:

1."\lambda =?; \nu=7.30*10^{15}s^{-1}"

We know; "\lambda=c/ \nu;" where "c=3*10^8 ms^{-1}" (speed of light)

"=3*10^8\/(7.30*10^{15})"

"=4.109*10^{-8}m."

"=41.096 nm."

2."\nu =?; E=3.8*10^{-10}J"

We know ; "E=h\nu" ; where "h=6.626*10^{-34} Js" (Planck's constant)

"\implies \nu=E/h"

"=3.8*10^{-10}\/(6.626*10^{-34})"

"=5.735*10^{23}s^{-1}" .